# Decoding Geometric Properties in Non-Random Data from First Information-Theoretic Principles

### v1

@@ 2.1 On the perturbation analysis of compressed data
Section $4$ of the Sup. Inf. presents empirical results that show our method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space. Our method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Section 2.2), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
@@
In particular, the Block Decomposition Method (BDM) [30] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes displayed in Figures $1$, $4$, $6$, $8$, and $9$ in Section $4$ of the Sup. Inf.. Additionally, as shown in Figure $6$, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise, whose impact on our message reconstruction method is investigated in Figure $5$. As shown in Figures $2$, $3$, and $7$, algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [31, 30].
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will demonstrate in Section 2.2, we have found sufficient conditions for which the method remains sound for any multidimensional space. The challenges to tackle this generalization is presented in Section $3$ of the Sup. Inf..
@@ 2.2 Zero-knowledge one-way communication
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this manner, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Section $5$ of the Sup. Inf. for further discussion on this and other mathematical phenomena. The proofs of all theorems can be found in Section $3$ of the Sup. Inf..
@@ 3 Discussion
As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Further discussion of these and other topics can be found in Section $5$ of the Sup. Inf..
@@ 4.2 Toward a general theory for zero-knowledge one-way communication
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [35, 36]. (See also Section $2$ in the Sup. Inf..) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem $3.10$ in the Sup. Inf.).
@@
As further discussed in Section $5$ of the Sup. Inf., we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.

### v2

@@ 2.1 On the perturbation analysis of compressed data
Section A.4 presents empirical results that show our method provides a robust approach to identifying the best way to decode a signal back to its original multidimensional space. Our method consists in sweeping over a large variety of possible dimensional configurations and measuring the resulting information content and algorithmic complexity of the candidate messages by taking the lowest algorithmic complexity configurations under the assumption that the original message is algorithmically not random (i.e., that the original message is compressible). As shown in [29] for graphs, this is because if the original message is random, then the likelihood that a rearrangement (i.e., a new partition) will lead to a low complexity configuration is very low. On the other hand, if the original message is sufficiently compressible (see Section 2.2), and therefore not random, most partitions will lead to configurations of greater algorithmic randomness.
@@
In particular, the Block Decomposition Method (BDM) [30] (combining Shannon entropy for long ranges and algorithmic probability for short ranges) was shown to be the complexity measure that was most sensitive and accurate in identifying the original message (image) across the complexity landscapes displayed in Figures $1$, $4$, $6$, $8$, and $9$ in Section A.4. Additionally, as shown in Figure $6$, BDM significantly outperforms Compression and Shannon Entropy in the face of additive noise, whose impact on our message reconstruction method is investigated in Figure $5$. As shown in Figures $2$, $3$, and $7$, algorithmic complexity (as estimated via BDM) remains invariant to low-algorithmic-complexity transformations that preserve the algorithmic content of the data (such as swapping all 0s and 1s or swapping all black and white pixels in an image), as well as some non-linear but algorithmic-information-content-preserving transformations [31, 30].
While the above discussed cases study the bidimensional spaces, Fig. LABEL:bone below shows that our method can be scaled up to tridimensional spaces. As we will demonstrate in Section 2.2, we have found sufficient conditions for which the method remains sound for any multidimensional space. The challenges to tackle this generalization is presented in Section A.2.
@@ 2.2 Zero-knowledge one-way communication
One of the main conditions for this to hold is that the algorithmic complexity of the object $y$ (with or without noise) in the wrong multidimensional space should increase (i.e., move the message toward algorithmic randomness) sufficiently fast in comparison to the algorithmic complexity of the message in a correct multidimensional space. To this end, one can assume a message should contain a large amount of redundancies, and thus should be highly compressible with respect to its linear signal stream length $\left|y\right|$, such as the case in which $\mathbf{K}\left(y\middle|\mathcal{S}\right)=\mathbf{O}\left(\log\log\left(\left|y\right|\right)\right)$ holds, where $\mathbf{K}\left(\cdot\right)$ denotes the algorithmic complexity. In this manner, most low-complexity perturbations will transform the object and/or partition by sufficiently moving it toward algorithmic randomness so that the conditions are satisfied. This condition resembles the ones for the noisy-channel coding theorem in classical information theory [20], although our method can be understood as being more general than the classical version because: it holds for zero-knowledge communication channels, while the classical version requires that the encoding-decoding schemes must be previously known by the emitter and receiver; and stochastic noise (as in the classical version) is a particular case of randomly generated algorithmic perturbations [26]. See also Section A.5 for further discussion on this and other mathematical phenomena. The proofs of all theorems can be found in Section 2.2.
@@ 3 Discussion
As shown in [24, 32], AID allows the investigation of causal analysis [28, 29] and solution of inverse problems [25] by taking advantage of a high convergence rate to the algorithmic probability brought along by BDM. In addition, although it may seem paradoxical at first glance—because one may initially think that the errors produced by the methods for approximating the algorithmic complexity would be inherited by the perturbation analysis phase itself—our theoretical and empirical results show that the perturbation analysis enables one to overcome the semi-computability limitation in algorithmic information theory, but only when, looking at the whole landscape of complexity values, one estimates with a sufficiently large amount of perturbations effected on the received signal. This is because the algorithmic complexities of the perturbed signals eventually become sufficiently larger than the complexity of the original message. Hence, a powerful enough method for approximating complexity, such as BDM, can eventually harness this divergence in order to enable the perturbation analysis to statistically distinguish the correct one (i.e., more compressible) from the wrong ones (i.e., more random). That is, an approximating method with enough computational resources can produce smaller errors than the difference of actual (although semi-computable) complexity values between the correct candidate and the perturbed ones, even though the exact complexity values remain ultimately unknown to the perturbation analyst (i.e., even though the analyst itself can never be sure it has enough computational resources in the first place). Further discussion of these and other topics can be found in Section A.5.
@@ 4.2 Toward a general theory for zero-knowledge one-way communication
In the general case (where the multidimensional spaces may be much more complex than the objects themselves), even in the ideal scenario in which we correctly infer the actual object sent by a completely unknown emitter, we are still prone to being wrong about the original dimensions’ configurations the emitter intended [35, 36]. (See also Section A.2 ) This poses a challenge to straightforward applications of the algorithmic information theory (AIT). In order to tackle such a challenge, we develop a method for multidimensional reconstruction of received low-dimensional signals in which one does not have access to knowledge about the encoding-decoding scheme or the original multidimensional space. Theorem 1 demonstrates the existence of sufficient conditions (or assumptions) that allow one to achieve an encoding/decoding-agnostic method for reconstructing the original multidimensional spaces. It gives rise to an agnostic method that does not need to assume that one a priori knows that the emitter agent’s multidimensional space resembles ours, is as complex as ours, or is isomorphic to the structures we currently deem fundamental in mathematics.
Theorem 1 (Theorem 11 in the Sup. Inf.).
@@
As further discussed in Section A.5, we believe the results in this article not only contribute with a general method for interpreting messages in zero-knowledge one-way communication channels, but also establish fundamental connections (and posit open problems) with zero-knowledge proofs and the algorithmic counterparts of the maximum likelihood method and the noisy-channel coding theorem.